OpenStudy (anonymous):

Solve each equation. (x + 3)(x - 3) = 16 - 2x2

OpenStudy (anonymous):

multiply left side things.. u will get x2-9 now x2-9=16-2x2 3x2=16+9 x2=25/3 x=+-5sqrt(1/3)

OpenStudy (yanasidlinskiy):

Simplifying (x + 3)(x + -3) = 16 + -2x^2 Reorder the terms: (3 + x)(x + -3) = 16 + -2x^2 Reorder the terms: (3 + x)(-3 + x) = 16 + -2x^2 Multiply (3 + x) * (-3 + x) (3(-3 + x) + x(-3 + x)) = 16 + -2x^2 ((-3 * 3 + x * 3) + x(-3 + x)) = 16 + -2x^2 ((-9 + 3x) + x(-3 + x)) = 16 + -2x^2 (-9 + 3x + (-3 * x + x * x)) = 16 + -2x^2 (-9 + 3x + (-3x + x^2)) = 16 + -2x^2 Combine like terms: 3x + -3x = 0 (-9 + 0 + x^2) = 16 + -2x^2 (-9 + x^2) = 16 + -2x^2 Solving -9 + x^2 = 16 + -2x^2 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x^2' to each side of the equation. -9 + x^2 + 2x^2 = 16 + -2x^2 + 2x^2 Combine like terms: x2 + 2x2 = 3x2 -9 + 3x2 = 16 + -2x2 + 2x2 Combine like terms: -2x2 + 2x2 = 0 -9 + 3x2 = 16 + 0 -9 + 3x2 = 16 Add '9' to each side of the equation. -9 + 9 + 3x2 = 16 + 9 Combine like terms: -9 + 9 = 0 0 + 3x2 = 16 + 9 3x2 = 16 + 9 Combine like terms: 16 + 9 = 25 3x2 = 25 Divide each side by '3'. x2 = 8.333333333 Simplifying x2 = 8.333333333 Take the square root of each side: x = {-2.886751346, 2.886751346}

OpenStudy (anonymous):

Solve each equation. 0.2x2 + 1.2 = 3.4

OpenStudy (yanasidlinskiy):

Not really sure about this one...But here's my: 0.2x2 + 1.2 = 3.4 divide by 0.2=2/10 {that is multiply by 10/2} x2+6=17 --> x2-11=0 (x-sqrt(11))(x+sqrt(11))=0

OpenStudy (anonymous):

I have one more is that okay?

OpenStudy (yanasidlinskiy):

Go ahead:)

OpenStudy (anonymous):

Solve each equation. 6x^2 + 1 = 13 and then Solve this equation 4 = x^2 -7

OpenStudy (yanasidlinskiy):

For the first one: Simplifying 6x2 + 1 = 13 Reorder the terms: 1 + 6x2 = 13 Solving 1 + 6x2 = 13 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + 6x2 = 13 + -1 Combine like terms: 1 + -1 = 0 0 + 6x2 = 13 + -1 6x2 = 13 + -1 Combine like terms: 13 + -1 = 12 6x2 = 12 Divide each side by '6'. x2 = 2 Simplifying x2 = 2 Take the square root of each side: x = {-1.414213562, 1.414213562}

OpenStudy (yanasidlinskiy):

I'm not sure about the second one... @paki

OpenStudy (paki):

it will be.... 4 = x^2 -7 x^2=4+7 x^2=11... now take the root on both sides... x=3.31.... @YanaSidlinskiy